OpenStudy (bekkah323):

x+4/x plus x/x+4 * equation below

OpenStudy (bekkah323):

\[\frac{ x+4 }{ x } + \frac{ x }{ x+4 }\]

OpenStudy (anonymous):

Try common denominators. \[\frac{ (x+4)(x+4) + (x)(x) }{ (x)(x+4) }\] just expand and simplify.

OpenStudy (bekkah323):

so would it be \[\frac{ 2x^2+8x+16 }{ x(x+4)}\]

OpenStudy (anonymous):

sorry about that watching some tennis lol.... yea that is right